OpenStudy (anonymous):

Solve each system by graphing; 2x-y=-2 x-3y=9 Thanks:)

OpenStudy (anonymous):

y=2(x+1)

OpenStudy (anonymous):

y=x-9/3

OpenStudy (anonymous):

how did you get those?:o

OpenStudy (anonymous):

im confused on the steps /:

OpenStudy (mysesshou):

2x-y=-2 x-3y=9 First one: 2x-y=-2 subtract 2x both sides: -y=-2-2x multiply both sides by -1 : y=2+2x y=2x+2 y intercept : (0,2) x intercept: set y = 0 0=2x+2 2x=-2 x=-1 (-1,0) Second one: x-3y=9 subtract x both sides: -3y=9-x divide both sides by -3 y=(9/(-3))-(-1/3)x y=(-3)+(1/3)x y=(1/3)x -3 y intercept : (0,-3) x intercept: set y = 0 0=(1/3)x -3 (1/3)x = 3 x=9 (3,0)

OpenStudy (mysesshou):

 oops. the last line I had typed should say (9,0)